Do everything
6A Smal tutle moves at a Speed o 134 e wr longs perrtFinde Speed 0s ond use n cm per second. Note tho 1 fun lon fortnight- IH days 23. A Can is traweling at Spe of 33.omls on an interstate highuay ushere ihe Speed timi s 15.0 milh. s the driver k Ceeding Speed imt? Youn answer. 5 The diometr of a Sphere is measuned to be 5.36 in. Find. a. The radiun 아 He Sphre in b. The Sunfate anea of tte Sphere in Squane cm, and The yotume ot he sphere in ubit m.
Answer #1

Answer(23).

75.0 mi/hr = 75*.447 = 33.25 m/s^2

1 mi/hr (mi hr-1) = 1.47 ft/s (ft s-1) = 1.61 km/hr (km hr-1) = 0.447 m/s (m s-1)

Answer(25).

Given diameter of sphere = 5.36 inches

= 5.36 * 2.54 centimetres (because 1 inch = 2.54 centimetres)

= 13.6144 centimetres


Radius of sphere = diameter/2

= 13.6144/2 centimetres

= 6.8072 centimetres


Surface area of sphere = 4*pi*radius*radius

= 4 * (22/7) * 6.80 * 6.80 square centimetrs

= 581.30 square centimetres


Volume of sphere = (4/3) * pi * radius * radius * radius

= (4/3) * (22/7) * 6.80 * 6.80 * 6.80 cubic centimetres

= 1317.619 cubic centimetres